OpenStudy (nightmarenight):

Is triangle ABC with vertices A(-1,4), B(3,1), and C(0,-3) a right triangle? Explain your answer using the slopes of the segments that form the sides of the triangle.

OpenStudy (anonymous):

Umm, did you try to graph it?

OpenStudy (andrewhaze):

that's what I was thinking too

OpenStudy (andrewhaze):

I mean do you have to graph it

OpenStudy (anonymous):

You can find the slope between any two points by using: \[ \frac{y_2-y_1}{x_2-x_1}=m \]

OpenStudy (anonymous):

You know that two slopes are perpendicular if \(m_1m_2=-1\).

OpenStudy (nightmarenight):

It's not a right triangle, I did graph it

OpenStudy (nightmarenight):

But I don't think graphing it was required

OpenStudy (andrewhaze):

but wouldn't you have to graph it to find out if it was a right triangle or not

OpenStudy (anonymous):

Can you find slope for the AB segment using the formula I wrote?

OpenStudy (nightmarenight):

I don't get slopes at all qq

OpenStudy (anonymous):

The slope is the difference of the y coordinates divided by the difference of the x coordinates

OpenStudy (andrewhaze):

^

OpenStudy (anonymous):

So, for AB, we have \[ \frac{A_y-B_y}{A_x-B_x} = \frac{4-1}{-1-3} \]

OpenStudy (nightmarenight):

Ohh, so it's -1+3+0?

OpenStudy (andrewhaze):

is that what you got

OpenStudy (nightmarenight):

\[\frac{ 3 }{ -4 }\]

OpenStudy (anonymous):

Okay, now can you find slope for AC and BC?

OpenStudy (nightmarenight):

1,7 and 3,4 @wio

OpenStudy (anonymous):

slope should be a fraction or whole number

OpenStudy (anonymous):

Anyway, you have to multiply each two slopes, and if any of them multiply to \(-1\), then that means they're perpendicular.

OpenStudy (nightmarenight):

I'm so lost;;;;

OpenStudy (anonymous):

So, for AC, we have \[ \frac{A_y-C_y}{A_x-C_x} = \frac{4-(-3)}{-1-0} \]